### v1

# Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

Abstract
Standard deep learning architectures such as convolutional neural networks and vision transformers often fail to generalize to previously unseen domains due to the implicit assumption that both source and target data are drawn from independent and identically distributed (i.i.d.) populations. In response, Domain Generalization techniques aim to enhance model robustness by simulating novel data distributions during training, typically through various augmentation or stylization strategies. However, these methods frequently suffer from limited control over the diversity of generated images and lack assurance that these images span distinct distributions. To address these challenges, we propose FDS, a novel strategy that employs diffusion models to synthesize samples from new domains by training on source distribution samples and performing domain mixing. By incorporating images that pose classification challenges to models trained on original samples, alongside the original dataset, we ensure the generation of a training set that spans a broad distribution spectrum. Our comprehensive evaluations demonstrate that this methodology sets new benchmarks in domain generalization performance across a range of challenging datasets, effectively managing diverse types of domain shifts. The implementation is available at: https://github.com/Mehrdad-Noori/FDS.git.
1 Introduction
Deep learning architectures, including Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have significantly advanced the field of computer vision, achieving state-of-art results in tasks like classification, semantic segmentation, and object detection. Despite these advancements, such models commonly operate under the simplistic assumption that training data (source domain) and post-deployment data (target domain) share identical distributions. This overlook of distributional shifts results in performance degradation when models are exposed to out-of-distribution (OOD) data [1, 2]. Domain adaptation (DA) [3, 4, 5] and Test-Time Adaptation (or Training) [6, 7, 8] strategies have been developed to mitigate this issue by adjusting models trained on the source domain to accommodate a predefined target domain. Nonetheless, these strategies are constrained by their dependence on accessible target domain data for adaptation, a prerequisite that is not always feasible in real-world applications. Furthermore, adapting models to each novel target domain entails considerable computational overhead, presenting a practical challenge to their widespread implementation.
Domain generalization (DG) [9] aims to solve the issue of domain shift by training models using data from one or more source domains, so they perform well out-of-the-box on new, unseen domains. Recently, various techniques have been developed to tackle this problem [10, 11], including domain aligning [12, 13, 14], meta-learning [15, 16], data augmentation [17, 18, 19], ensemble learning[20], applying self-supervised learning techniques [21, 22], and adding regularization methods [23, 24]. These strategies are designed to make models more adaptable and capable of handling data that they were not explicitly trained on, making them more useful in real-world situations where the exact nature of future data cannot be predicted. Among these techniques, a notable category focuses on synthesizing samples from different distributions to mimic target distributions. This is achieved through strategies like image transformation [17, 25], style transfer [26, 27], learnable augmentation networks [28, 29, 30], and feature-level stylization [20, 31]. However, many of these methods face challenges in controlling the synthesis process, often resulting in limited diversity where primarily only textures are altered.
Figure 1: Bridging Domain Gaps with FDS: Comprehensive Distribution Coverage Across Domains.
In this study, we introduce an innovative approach using diffusion model to address the challenge of domain generalization. Known for their exceptional ability to grasp intricate distributions and semantics, diffusion models excel at producing high-quality, realistic samples [32, 33, 34, 35]. We exploit this strength by training a diffusion model that is conditioned on various classes and domains present in the training dataset, aiming to master the distribution of source domains. Through the process of domain mixing during generation, we create images that appear to originate from new, pseudo-target domains. To ensure the development of a robust classifier, we initially select generated samples that are difficult for a model trained solely on the original source domains to classify. Subsequently, we train the model using a combination of these challenging images and the original dataset. In this manner, we ensure that our model is fed with images of the widest possible diversity, thereby significantly enhancing its ability to generalize across unseen domains. Our contributions can be summarized as follows:
  • We introduce a novel DG approach that leverages diffusion models conditioned on different domains and classes to generate samples from novel, pseudo-target domains through domain interpolation. This approach increases the diversity and realism of the generated images;
  • We propose an innovative strategy for selecting images that pose a challenge to a classifier trained on original images, ensuring the diversity of the final sample set. We then train a classifier on this enriched dataset, comprising both challenging generated images and original images, to enhance its generalization capabilities;
  • We conduct extensive experiments across various benchmarks to validate the effectiveness of our method. These experiments demonstrate FDS’s ability to significantly improve the robustness and generalization of models across a wide range of unseen domains, achieving state-of-art performance on these datasets.
The rest of this paper is structured as follows. The next section reviews previous work on domain generalization and recent uses of diffusion models. Section 4 defines the DG problem we focus on and introduces our proposed method. Section 5 presents the experimental setup employed to validate our method. In Section 6, we share the results and findings of our experiments, showing how our method improves upon previous work. We wrap up the paper in Section 7, discussing our main results and ideas for what could be studied next.
2 Related Works
2.1 Domain Generalization (DG)
Domain Generalization, a concept first introduced by Blanchard et al. in 2011 [9], has seen growing interest in the field of computer vision. This interest has spurred the development of a broad spectrum of methods aimed at enabling models to generalize across unseen domains. These methods can be categorized into several key approaches, each targeting the DG problem from a unique angle.
Domain alignment.  Strategies based on domain alignment focus on minimizing the disparity between source domains by learning representations that are invariant across these domains. Various techniques, such as moment matching [36], discriminant analysis [12], minimizing maximum mean discrepancy (MMD)[37], employing contrastive loss[38], and domain-adversarial learning [39], are employed to achieve this goal.
Meta-learning.  Approaches using meta-learning address DG [15, 16] by considering a bi-level optimization problem, where the model is fine-tuned on meta-source domains to minimize the error on a meta-target domain. This is often achieved through episodic training that may involve updating all [15] or a subset [16] of the network’s parameters.
Data augmentation.  A key strategy in DG, data augmentation aims to enhance model robustness against domain shifts encountered during deployment. This objective is pursued through a variety of techniques, such as learnable augmentation, off-the-shelf style transfer, and augmentation at the feature level. Learnable augmentation models utilizes networks to create images from training data, ensuring their distribution diverges from that of the source domains [28, 30, 29]. Meanwhile, off-the-shelf style transfer based methods seek to transform the appearance of images from one domain to another or modify their stylistic elements, often through Adaptive Instance Normalization (AdaIN)[40, 26]. Unlike most augmentation approaches that modify pixel values, some propose altering features directly, a technique inspired by the finding that CNN features encapsulate style information[41, 20] in their statistics.
Ensemble learning.  Techniques based on this approach focus on improving robustness to out-of-distribution (OOD) data by developing multiple models tailored to specific domains [20]. The collective predictions of these models, weighted by either the resemblance of a target sample to source domains or the confidence levels of the models, offer a nuanced assessment for target domain samples [42].
Self-supervised learning.  Conversely, self-supervised learning methods concentrate on fostering domain-agnostic representations through the pre-training of models on unsupervised tasks, such as puzzle solving [21, 43], image reconstruction with autoencoders [44], or engaging in contrastive learning [45].
Disentangled representation learning  Methods in this category try to segregate features into those unique to specific domains and those common across all domains [46, 47]. This segregation is facilitated by domain-specific masks that highlight the features most pertinent to images from the target domain [47].
Regularization  Lastly, regularization methods in DG are designed to refine the training process, encouraging the development of features that are universally applicable across domains [24, 48, 49, 50]. These methods typically build on the empirical risk minimization (ERM) framework [51], incorporating additional objectives such as distillation [48, 50], stochastic weight averaging [24], or distributionally robust optimization (DRO) [49] to promote generalization.
2.2 Diffusion Models
Diffusion models have recently surpassed Generative Adversarial Networks (GANs) as the leading technique for image synthesis. Innovations like denoising diffusion probabilistic models (DDPMs) [33] and denoising diffusion implicit models (DDIMs) [35] have significantly sped up the image generation process. Rombach et al. [34] introduced latent diffusion models (LDMs), also known as stable diffusion models, which enhance both training and inference efficiency while facilitating text-to-image and image-to-image conversions. Following this advancement, various enhancements and extensions have been developed, including stable diffusion XL (SDXL) [52] and ControlNet [53]. These models introduce the capability to guide the generation process using additional inputs, such as depth or semantic information, allowing for more controlled and versatile image creation.
Despite the impressive capabilities of diffusion models in generating images, their application in domain generalization has been minimally explored. Building on the concept of using style transfer networks, Yu et al. have applied diffusion models to adapt the style of target images onto source images [54], though this approach incurs additional computational costs during deployment. In a recent development, Niemeijer et al. introduced DIDEX [55], which leverages textual inputs to alter the style of input images, thereby generating a pseudo-target dataset that enhances model performance through adaptation to these pseudo-target samples within the unsupervised domain adaptation (UDA) framework. However, this method is constrained by the need for human intervention in selecting text inputs that align with the anticipated characteristics of the target domain. Diverging from the strategy of stylizing input images, our method adopts a unique approach by directing the diffusion process to produce new images from novel, diverse distributions while preserving the semantic integrity of the source samples.
3 Theoretical Motivation
Our classifier, represented by $f$ with parameters $\theta$, aims to craft a unified model from $n$ source domains $\{\mathcal{D}^{1},\ldots,\mathcal{D}^{n}\}$ that adapts to a novel target domain $\mathcal{D}^{\mathcal{T}}$. Within any domain $\mathcal{D}$, we measure classification loss by:
$\mathcal{L}_{\mathcal{D}}(\theta)\,=\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\,\big{[}\ell(f(x;\theta),y)\big{]},$ (1)
where $x$ and $y$ denote the input and its corresponding label, respectively, and $\ell(f(x;\theta),y)$ is the cross entropy loss.
Empirical Risk Minimization (ERM) [56] forms the foundation for training our models, aiming to reduce the mean loss across training domains:
$\min_{\theta}\ \sum_{i}\frac{1}{|\mathcal{D}^{i}|}\sum_{k=1}^{|\mathcal{D}^{i}|}\ell(f(x_{i}^{k};\theta),y_{i}^{k})$ (2)
Here, $|\mathcal{D}^{i}|$ counts the data points in domain $i$, with $x_{i}^{k}$ as the $k$-th point and $y_{i}^{k}$ as its label. However, the accuracy of ERM-trained models drops when data shifts occur across domains due to inadequate OOD generalization. To enhance ERM, Chapelle et al. introduced Vicinal Risk Minimization (VRM) [57], which substitutes point-wise estimates with density estimation in the vicinity distribution around each observation within each domain. Practical implementation often involves data augmentation to introduce synthetic samples from these density estimates. Traditional augmentation processes a single data point $\tilde{x}_{i}^{k}$ from domain $i$ to yield $\tilde{x}_{i}^{k}=g(\tilde{x}_{i}^{k})$, where $g(\cdot)$ denotes a basic transformation and $\tilde{x}_{i}^{k}$ is the modified data point.
Despite the potential of VRM to boost performance on OOD samples, it does not completely bridge domain gaps. The root cause lies in the inability of ERM methods to anticipate shifts in data distribution, which simple augmentation within domains does not address. Hence, domain generalization requires more robust transformations to extend the model’s applicability beyond training domains. Muller et al. [58] have demonstrated that a wider range of transformations outperforms standard methods. However, Aminbeidokhti et al. [59] advise that aggressive augmentations could distort the essential characteristics of images, pointing to the necessity of a mechanism to filter out extreme alterations. The emergence of diffusion models, proficient in reproducing diverse data distributions, facilitates such advanced sampling approaches. Our goal, therefore, is to generate images that span across domain gaps, lessen the variability between data distributions, and introduce a system to exclude trivial or excessive modifications.
4 Method
Figure 2: Overview of the proposed architecture for FDS. (left) Multi-Source Training of Diffusion Model Conditioned on Class and Domain of the training images. (right) generating pseudo-novel image set using proposed interpolation and filtering mechanism of FDS.
To advance the generalization ability of our classifier, we aim to utilize generated image samples that traverse domain gaps yet retain class semantics. This objective is realized through a three-step methodology:
  • Step 1: We begin by training an image generator that masters the class-specific distributions within each domain, enabling the synthesis of class-consistent samples within those domains.
  • Step 2: Then, we introduce a strategy for generating inter-domain images to span the domain gaps effectively.
  • Step 3: Last, we filter overly simplistic samples from synthetic inter-domain images and train the final model on this refined set alongside original domain images for enhanced OOD generalization.
In the follow subsections, we present each of these steps in details.
4.1 Image Generator
Diffusion Models are designed to approximate the data distribution $p(x)$ by reversing a predefined Markov Chain of $T$ steps, effectively denoising a sample in stages. These stages are modeled as a sequence of denoising autoencoder applications $e_{\theta}(x_{t},t)$, for $t=1\ldots T$, which gradually restore their input $x_{t}$. This iterative restoration is formalized by the following objective:
$\mathcal{L}_{DM}\,=\,\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (3)
[…]
$\mathcal{L}_{LDM}\,=\,\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t)\right\|^{2}\right].$ (4)
Diffusion models can also handle conditional distributions $p(z|c)$. To enable this conditioning, building upon the work of Rombach et al. [34], we utilize a condition encoder $\tau_{\theta}$ that maps $c$ onto an intermediate representational space $\tau(c)\in\mathbb{R}^{M\times d_{\tau}}$. Following Stable Diffusion, we employ the CLIP-tokenizer and implement $\tau$ as a transformer to infer a latent code. This representation is subsequently integrated into the UNet using a cross-attention mechanism, culminating in our final enhanced objective:
$\mathcal{L}_{LDM}^{\mathrm{cond}}\,=\,\mathbb{E}_{\mathcal{E}(x),c,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau(c))\right\|^{2}\right],$ (5)
This approach enables a single diffusion model to understand and generate images across different domains for each class, using text-based conditions (prompts). By dynamically adjusting conditions, the model efficiently learns varied representations without needing multiple models for each domain. Thus, we employ a textual template $c(x)$, denoting “[y], [D]” where y is the class label and D is the domain name of the input $x$, and proceed to train our diffusion model using this template across all images from the source domains. Upon completing training, we possess the capability to create a new sample for class $y^{k}$, belonging to the set $\{y^{1},\ldots,y^{m}\}$, within domain $\mathcal{D}^{i}$. This is achieved by decoding a denoised representation after $t$ timesteps, $\tilde{x}_{t}^{i,k}=\tilde{\mathcal{E}}(\Phi_{t}(\mathcal{D}^{i},y^{k}))$, where $\Phi_{t}(\mathcal{D}^{i},y^{k})$ is the denoised representation that originates from random Gaussian noise conditioned on $\mathcal{D}^{i}$ and $y^{k}$.
4.2 Domain Mixing
As illustrated in 1, In order to leverage image generator to synthesize images from new unseen distributions we propose two mixing strategies:
4.2.1 Noise Level Interpolation.
Consider our dataset comprising $n$ source domains $\{\mathcal{D}^{1},\ldots,\mathcal{D}^{n}\}$ and target classes $y^{k}$ from the set $\{y^{1},\ldots,y^{m}\}$. Utilizing a trained image generator that initiates with random Gaussian noise, we denoise this input over $t$ steps to produce a synthetic, denoised representation $\tilde{z}_{t}^{i,k}=\Phi_{t}(\mathcal{D}^{i},y^{k})$ indicative of domain $\mathcal{D}^{i}$ and class $y^{k}$. To synthesize a sample that merges the characteristics of domains $\mathcal{D}^{i}$ and $\mathcal{D}^{j}$ for class $y^{k}$, we employ a single diffusion model, conditioned dynamically to capture the essence of both domains. This process aims to generate a sample that embodies the transitional features between these domains, effectively bridging the domain gap.
The model begins its process from the same initial random Gaussian noise, adapting its denoising trajectory under two distinct conditions, $\tau(\mathcal{D}^{i},y^{k})$ and $\tau(\mathcal{D}^{j},y^{k})$, up to a specific timestep $T$. This dual-conditioned approach ensures that the evolving representation up to $T$ incorporates influences from both domains, guided by the respective conditional inputs.
From timestep $T$ onwards, until the final representation is formed, the model blends the outputs from these dual paths at each step. Specifically, for each step $t>T$, we form a mixed representation $\tilde{z}_{t}$ as:
$\tilde{z}_{t}\,=\,\alpha\,\Phi_{t}(\mathcal{D}^{i},y^{k})\,+\,(1\!-\!\alpha)\,\Phi_{t}(\mathcal{D}^{j},y^{k}),$ (6)
where $\alpha$ is a predefined mixing coefficient that dictates the blend of domain characteristics in the output. This combined representation $\tilde{z}_{t}$ is then used as the basis for the model’s next denoising step, integrating features from both $\mathcal{D}^{i}$ and $\mathcal{D}^{j}$ for class $y^{k}$. Through this iterative mixing, the model ensures a gradual and cohesive fusion of domain-specific attributes, leading to a synthesized sample that seamlessly spans the gap between the domains for class $y^{k}$.
4.2.2 Condition Level Interpolation.
We also propose Condition Level Interpolation to generate images that effectively bridge domain gaps. This technique relies on manipulating the conditions fed into our diffusion model to guide the synthesis of new samples. Specifically, for a target class $y^{k}$ and two distinct domains, $\mathcal{D}^{i}$ and $\mathcal{D}^{j}$, we employ our encoder $\tau(c)$ to create separate condition representations for each domain-class pair: $(y^{k},\mathcal{D}^{i})$ and $(y^{k},\mathcal{D}^{j})$.
The core of this strategy involves blending these condition representations using a mix coefficient $\alpha$, leading to a unified condition:
$c_{\mathrm{mixed}}\,=\,\alpha\,\tau(c_{y^{k}},\mathcal{D}^{i})\,+\,(1\!-\!\alpha)\,\tau(c_{y^{k}},\mathcal{D}^{j}).$ (7)
This mixed condition $c_{\text{mixed}}$ then orchestrates the generation process from the initial step, ensuring that the diffusion model is consistently influenced by attributes from both domains. By initiating this conditioned blending from the beginning of the diffusion process, we ensure a harmonious integration of domain characteristics throughout the generation of the synthetic image.
4.3 Filtering Mechanism
[…]
The utility of these synthetic samples in improving model generalization varies, prompting an entropy-based evaluation to identify those with the greatest potential. High entropy scores, indicating prediction uncertainty by a classifier $h(x)$ trained on the original dataset, suggest that such samples may come from previously unseen distributions. This characteristic posits these high-entropy samples as prime candidates for training, hypothesized to challenge the classifier significantly and aid in covering the domain gaps. Further refining this selection, we only include samples correctly predicted as their target class by $h(x)$, ensuring the exclusion of samples that have lost semantic integrity during the diffusion process.
Let $\mathcal{C}_{i,j}^{k}$ be the subset of samples in $S_{i,j}^{k}$ with are correctly classified by $h(x)$:
$\mathcal{C}_{i,j}^{k}\,=\,\big{\{}\tilde{x}_{i,j}^{k,(r)}\,|\,h(\tilde{x}_{i,j}^{k,(r)})=y^{k}\big{\}}.$ (8)
We choose from $\mathcal{C}_{i,j}^{k}$ the $N_{L}$ samples with highest entropy (the entropy of a $k$-class discrete probability distribution $p$ is given by $\mathcal{H}(p)=-\sum_{k}p_{k}\log p_{k}$) to form a set of selected samples $\tilde{\mathcal{D}}_{i,j}^{k}$. Last, we combine the synthetic samples, created for each class and domain pair, to the original dataset $\mathcal{O}$ to obtain the final augmented training set
[…]
Training the final classifier on $\mathcal{A}$ not only enriches the dataset but also ensures robust model generalization across diverse domain landscapes.
5 Experimental Setup
Table 1: Comparison to the state-of-art on three benchmarks, reporting the mean and standard deviation across three runs.The best results are in bold fonts.
Method Category PACS VLCS OfficeHome Avg. ERM [51] Baseline 85.5 $\pm$0.2 77.5 $\pm$0.4 66.5 $\pm$0.3 76.50 $\pm$0.3 Mixstyle [20] Data Augmentation 85.2 $\pm$0.3 77.9 $\pm$0.5 60.4 $\pm$0.3 74.50 $\pm$0.4 RSC [23] 85.2 $\pm$0.9 77.1 $\pm$0.5 65.5 $\pm$0.9 75.93 $\pm$0.8 Mixup [60] 84.6 $\pm$0.6 77.4 $\pm$0.6 68.1 $\pm$0.3 76.70 $\pm$0.5 SagNet [61] 86.3 $\pm$0.2 77.8 $\pm$0.5 68.1 $\pm$0.1 77.40 $\pm$0.3 DCAug [59] 86.1 $\pm$0.7 78.6 $\pm$0.4 68.3 $\pm$0.4 77.66 $\pm$0.5 IRM [62] Domain-Invariant 83.5 $\pm$0.8 78.5 $\pm$0.5 64.3 $\pm$2.2 75.43 $\pm$0.5 GroupDRO [49] 84.4 $\pm$0.8 76.7 $\pm$0.6 66.0 $\pm$0.7 75.70 $\pm$0.7 DANN [63] 83.6 $\pm$0.4 78.6 $\pm$0.4 65.9 $\pm$0.6 76.03 $\pm$0.5 MMD [64] 84.6 $\pm$0.5 77.5 $\pm$0.9 66.3 $\pm$0.1 76.13 $\pm$0.5 SelfReg [45] 86.5 $\pm$0.3 77.5 $\pm$0.0 69.4 $\pm$0.2 77.80 $\pm$0.2 CORAL [65] 86.2 $\pm$0.3 78.8 $\pm$0.6 68.7 $\pm$0.3 77.90 $\pm$0.4 mDSDI [66] 86.2 $\pm$0.2 79.0 $\pm$0.3 69.2 $\pm$0.4 78.13 $\pm$0.3 ERM(reproduced[51] 84.30 $\pm$1.1 76.23 $\pm$1.1 64.60 $\pm$1.1 75.04 $\pm$0.8 ERM + FDS (ours) 8 7.38 $\pm$1.6 79.46 $\pm$1.6 69.17 $\pm$1.1 78.67 $\pm$1.4 SWAD [24] Ensemble Learning 88.1 $\pm$0.1 79.1 $\pm$0.1 70.6 $\pm$0.2 79.27 $\pm$0.1 TeachDCAug [59] 88.4 $\pm$0.2 78.8 $\pm$0.4 70.4 $\pm$0.2 79.20 $\pm$0.3 DNA [67] 88.4 $\pm$0.1 79.0 $\pm$0.1 71.2 $\pm$0.1 79.53 $\pm$0.1 DIWA [68] 88.8 $\pm$0.4 79.1 $\pm$0.2 71.0 $\pm$0.1 79.63 $\pm$0.2 SWAD (reproduced[24] 88.11 $\pm$0.4 78.87 $\pm$0.5 70.34 $\pm$0.4 79.11 $\pm$0.4 SWAD + FDS (ours) 90.50 $\pm$0.3 79.73 $\pm$0.5 73.51 $\pm$0.4 81.25 $\pm$0.4
5.1 Dataset
Following Gulrajani and Lopez-Paz [51], we compare our proposed approach to the current state-of-art using three challenging datasets with different characteristics: PACS [46], VLCS [69], and OfficeHome [70].
The PACS dataset [46] has a total of 9,991 photos divided into four distinct domains, $d\in$Art, Cartoon, Photo, Sketch} and seven distinct classes. The second dataset, VLCS, comprises 10,729 photos from four separate domains, $d\in$Caltech101, LabelMe , SUN09, VOC2007} and five different classes. The third dataset, OfficeHome, includes a total of 15,588 photos taken from four domains, $d\in$Art, Clipart, Product, Real}, and 65 classes.
5.2 Implementation Details
To ensure a fair comparison, we adopt the DomainBed framework [51], a comprehensive benchmark that encompasses prominent domain generalization (DG) methodologies under a uniform evaluation protocol. Following this framework, we employ a leave-one-out strategy for DG dataset assessment where one domain serves as the test set while the others form the training set. A subset of the training data, constituting 20%, is designated as the validation set [^1] The model demonstrating peak accuracy on this validation set is selected as optimal. This model undergoes evaluation on the test domain, providing the unseen domain accuracy.. The aggregate result for each dataset represents the mean accuracy derived from varying the test domain. To ensure reliability, experiments are replicated three times, each with a unique seed.
Moreover, to rigorously test our method, we base our classifiers on the SWAD technique [51], a stable ensemble approach surpassing several recent DG methods. Our method’s efficacy is showcased without extensive hyperparameter tuning, adhering to SWAD’s recommended settings.
For image synthesis, the original Stable Diffusion framework is utilized with $\mathrm{DDIM}\!=\!50$ steps. The PACS and VLCS datasets prompt the generation of $N\!=\!32,000$ samples per class, whereas OfficeHome, with its 65 classes, necessitates $N\!=\!16,000$ samples. Image generation spans an interpolation range of $\alpha\!\in\![0.3,0.7]$ and a Noise Level Interpolation range of $\mathcal{T}\!\in\![20,45]$. This diversified parameter selection, rather than optimizing hyperparameters per dataset, acknowledges each domain’s unique shift. Our filtering mechanism then identifies the most informative images from the generated pool. The selection of an optimal dataset size, $N_{L}$, underwent extensive exploration, detailed further in Section 6.2.
6 Results
Table 2: Comparative Analysis of FDS Component Effects Across PACS Dataset Domains
Module Accuracy (%) Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 + Image Generation (No Interpol.) 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 $\pm$0.4 + Image Interpolation 91.38 $\pm$0.2 85.20 $\pm$0.6 97.73 $\pm$0.1 84.27 $\pm$0.9 89.65 $\pm$0.4 + Filtering Mechanism 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
We first compare the accuracy of our approach against state-of-art DG methods, across four benchmarks. We then present a detailed analysis investigating several key aspects of our approach, including the effectiveness of each proposed component, its regularization capabilities, and a comparison of different mixing strategies.
6.1 Comparison with the state of the art
In Table 1, we compare our approach with recent methods for domain generalization as outlined in the DomainBed framework [51]. Our method, when added to the standard ERM classifier, shows an improved accuracy of 3.1% on the PACS dataset using the ResNet-50 model. On the VLCS dataset, our method sees a 3.23% increase in accuracy over the standard ERM. For the OfficeHome dataset, our method outperforms the ERM baseline by 4.57%. To test the strength of our method, we also applied it to SWAD, which is a strong baseline for domain generalization. Here, our method improves the performance on the PACS dataset by 2.39%, on the VLCS dataset by 0.95%, and on the OfficeHome dataset by 3.17%. Detailed results of accuracy on each domain of all datasets can be found in Supplemental Material.
6.2 Further Analysis
Ablation Study on different Components.  To evaluate the impact of each component on out-of-distribution (OOD) generalization, we systematically introduce each module and observe the enhancements. As detailed in Table 2 first incorporating domain-specific synthetic samples, generated without interpolation, yields a $0.91\%$ gain over the baseline. This increment validates our diffusion model’s proficiency in capturing and replicating the class-specific distributions within domains, thus refining OOD performance through enhanced density estimation near original samples. Subsequently, integrating images from pseudo-novel domains via our interpolation strategy leads to a further $0.5\%$ enhancement, underscoring the mechanism’s effectiveness in connecting distinct domains. Lastly, applying our entropy-based filtering to eliminate overly simplistic images results in a significant $2.4\%$ improvement against the SWAD model, a robust benchmark. These outcomes collectively underscore the efficacy of our approach in fortifying OOD generalization.
Scaling Law of Data Size.  Next, we assess the impact of the selected data size, denoted as $N_{L}$, on the augmented dataset’s efficacy. We consider the PACS dataset for this analysis, where the average number of images per class is 570. We systematically explore varying scales relative to this average class size, aiming to discern the optimal dataset size for enhancing out-of-domain (OOD) generalization. Our findings, detailed in Figure 4, reveal an initial improvement in OOD generalization with increased data size. However, excessively enlarging the dataset size begins to diminish the benefits of our filtering mechanism, as it incorporates a broader array of samples, including those that are overly simplistic and not conducive to model improvement.
Filtering Mechanism.  One might consider that the enhanced accuracy observed with our filtering mechanism might not solely be its merit but rather the result of constraining the sample size for a balanced final training set. To investigate this hypothesis, Figure 4 contrasts the outcome of selecting $N_{L}$ samples at random from all generated images per class against employing our entropy-based filtering strategy. The consistent improvement in out-of-domain (OOD) generalization across all domains, facilitated by our filtering approach, underscores its effectiveness.
Additionally, to examine the impact of excluding samples that deviate semantically which will be identified through misclassification by a classifier trained solely on the original dataset, we contrasted the accuracy between selections purely based on entropy and those refined this way in Table 3. This comparison highlights the significance and efficiency of our filtering strategy in preserving semantic integrity.
Table 3: Impact of Filtering Strategy Components on PACS Dataset Accuracy
Method Accuracy (%) PACS VLCS OfficeHome Avg. Baseline (SWAD [24]) 88.11 $\pm$0.4 78.87 $\pm$0.5 70.34 $\pm$0.4 79.11 $\pm$0.4 Filtered Based on Entropy 90.50 $\pm$0.4 79.33 $\pm$0.8 71.97 $\pm$0.3 80.60 $\pm$0.5 + Reject Extreme Transform. 90.50 $\pm$0.3 79.73 $\pm$0.5 73.51 $\pm$0.5 81.25 $\pm$0.4
Figure 3: Accuracy Comparison on the PACS Dataset: Random Selection vs. Proposed Filtering Strategy of FDS
Figure 4: Impact of varying scales of sample size $N_{L}$ relative to the average number of images per class on Accuracy in the PACS Dataset Analysis.
Mixing Strategies.  Our ablation study, presented in Table 4, reveals that both Noise Level Interpolation and Condition Level Interpolation significantly outperform the baseline, yet their effectiveness varies with the dataset’s attributes and the nature of the domain shift. Noise Level Interpolation is optimal for minimal domain shifts, focusing on adjusting the noise aspect to bridge domain gaps. However, it falls short in scenarios with substantial domain differences, such as the transition to Cartoon or Sketch, where the source and target domains diverge significantly. In these cases, Condition Level Interpolation proves more advantageous, offering a robust mechanism for navigating complex domain shifts by manipulating higher-level semantic representations. This strategy is particularly effective when conventional noise adjustments are insufficient for achieving meaningful domain traversal.
Table 4: Out-of-Distribution Generalization Performance Using Various Interpolation Strategies of FDS
Strategy Accuracy (%) Art Cartoon Photo Sketch Avg. Baseline (SWAD [24]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 Noise Level Interpol. 91.95 $\pm$0.3 83.23 $\pm$0.1 97.56 $\pm$0.1 85.85 $\pm$0.3 89.65 $\pm$0.2 Condition Level Interpol. 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
6.2.1 In-Domain Regularization Effect.
In this section, we study the impact of incorporating synthetic images produced by our method on the accuracy of the network when tested on the same domains on which it was trained. As indicated in Table 5, FDS approach surpasses the baseline accuracy within this context. This affirms the value of our method, not only in out-of-distribution (OOD) conditions, but also in standard in-domain validation, aligning with the principles of Vicinal Risk Minimization (VRM) [57]. Hence, our method may also be viewed as a regularization strategy, suitable for a broad spectrum of applications.
Table 5: Effect of FDS on In-Domain Accuracy: PACS Dataset Analysis
Method Source Domains C, S, P A, P, S A, C, S A, C, P Avg. ERM [51] 97.72 $\pm$0.2 97.72 $\pm$0.2 96.54 $\pm$0.2 97.46 $\pm$0.4 97.36 $\pm$0.2 ERM + FDS (ours) 98.14 $\pm$0.4 97.86 $\pm$0.5 97.44 $\pm$0.2 98.04 $\pm$0.1 97.87 $\pm$0.2 SWAD [24] 98.23 $\pm$0.2 97.95 $\pm$0.4 97.30 $\pm$0.1 98.40 $\pm$0.2 97.97 $\pm$0.2 SWAD + FDS (ours) 98.43 $\pm$0.4 98.12 $\pm$0.1 97.90 $\pm$0.3 98.72 $\pm$0.2 98.29 $\pm$0.1
6.3 Stability Analysis
In this section, we demonstrate the performance of our model across different stages of training within two domains of the PACS dataset, depicted in Figure 5. It is important to note that these test accuracies were not used in the selection of the best-performing model mentioned in earlier sections and all of our experiments follow leave-one-out settings suggested by DomainBed protocol. A notable issue with domain shift, as evidenced by our findings, is the substantial and irregular fluctuations in test accuracy during training. This typically necessitates a comprehensive exploration of hyperparameter settings. For context, as reported in Table 6, training an ERM model within the DomainBed framework required 4,142 experiments to determine the most effective hyperparameters. In contrast, our approach yielded more consistent test accuracies over time, removing the need for such exhaustive hyperparameter searches. Consequently, we were able to achieve and report our results using a single, default set of hyperparameters. Details of hyperparameter search space for each framework is included in supplemental materials.
Figure 5: Test Accuracy Across Training Intervals: Comparison between ERM (Top Row) vs. FDS (Bottom Row) in Art and Sketch Domains of PACS
Framework #Experiments DomainBed [51] 4142 SWAD [24] 396 Ours 1
Table 6: Size of HP search space in different frameworks
7 Conclusion
This work presented FDS, a domain generalization (DG) technique that leverages diffusion models for domain mixing, generating a diverse set of images to bridge the domain gap between source domains distribution. We also proposed an entropy-based filtering strategy enriches the pseudo-novel generated set with images that test the limits of classifiers trained on original data, thereby boosting generalization. Our extensive experiments across multiple benchmarks demonstrate that our method not only surpasses existing DG techniques but also sets new records for accuracy. Our analysis indicate that our approach contributes to more stable training processes when confronted with domain shifts and serves effectively as a regularization method in in-domain contexts. Notably, our technique consistently enhances performance across diverse scenarios, from realistic photos to sketches, without necessitating extensive hyperparameter tuning commonly associated with other domain generalization strategies. Moreover, while we exploited our trained diffusion model for covering the domain gap, more sophisticated techniques could be considered. For instance, future work could investigate the idea of generating images by extrapolate in domain space to explore pseudo-novel distributions.
Appendix A Implementation
Our proposed FDS method is built using the Python language and the PyTorch framework. We utlized four NVIDIA A100 GPUs for all our experiments. For initializing our models, we utilize the original Stable Diffusion version 1.5 as our initial weight  [34]. The key hyperparameter configurations employed for training these diffusion models and generating new domains are detailed in Tables 7 and 8, respectively.
Furthermore, for classifier training, we adhere to the methodologies and parameter settings described by Cha et al. [24], ensuring consistency and reproducibility in our experimental setup. The original implementation and instructions for reproducing our results are accessible via our GitHub repository.
Appendix B Additional Ablation
Selection/Filtering. In this section, we provide visual examples to show the efficacy of our synthetic sample selection and filtering mechanism. As mentioned in the method section, this mechanism is intricately designed to scrutinize the generated images through two lenses: the alignment of the predicted class with the intended label, and the entropy indicating the prediction’s uncertainty.
The Figures 678 showcase a set of images generated from interpolations between two domains. Specifically, the diffusion model is trained on “art”, “sketch”, and “photo” of the PACS dataset, and the selected images, demonstrated in the first two rows, exemplify successful blends of domain characteristics, embodying a balanced mixture that enriches the training data with novel, domain-bridging examples. These images were chosen based on their ability to meet our criteria: correct class prediction aligned with high entropy scores. The third and fourth rows highlight the filtering aspect of our mechanism, displaying images not selected due to class mismatches and low entropy, respectively. This visual demonstration underlines the pivotal role of our selection/filtering process in refining the synthetic dataset, ensuring only the most challenging and domain-representative samples are utilized for model training. Through this approach, we aim to significantly bolster the model’s capacity to generalize across diverse visual domains.
Inter-domain Transition. In this section, we demonstrate the model’s ability to navigate between distinct visual domains, a capability enabled by adjusting the mix coefficient $\alpha$. Trained on multiple source domains, our model can generate images that blend the unique attributes of each source domain. By varying $\alpha$ from 0.0 to 1.0, we enable smooth transitions between two source domains, where $\alpha=0.0$ and $\alpha=1.0$ correspond to generating pure images of the first and second domain, respectively. As an example, we illustrated this ability for our model trained on the PACS sources’ “art”, “sketch”, and “photo”. These domain transitions are illustrated in the figures, showcasing transitions from “photo” to “art” domain in Figure 9, “sketch” to “art” domain in Figure 10, and “sketch” to “photo” domain in Figure 11, respectively. The examples provided highlight the effectiveness of our interpolation method in producing images that incorporate the distinctive features of the mixed domains, thus affirming the model’s capability to generate novel and coherent visual content that bridges the attributes of its training domains. Note that in all of our generation experiments, we constrained $\alpha$ to the range of 0.3 to 0.7 to ensure the generated images optimally embody the characteristics of the two mixing domains, as detailed in Table 8.
Table 7: Hyperparameter Configuration for Training Diffusion Models.
Config Value
Number of GPUs 4
Learning rate 1e-4
Learning rate scheduler LambdaLinear
Batch size 96 (24 per GPU)
Precision FP16
Max training steps 10000
Denosing timesteps 1000
Sampler DDPM [33]
Autoencoder input size 256 x 256 x 3
Latent diffusion input size 32 x 32 x 4
Table 8: Hyperparameter Configuration for Generating New Domains.
Config Value
Sampler DDIM [35]
Denosing timesteps 50
Classifier-free guidance (CFG) Randomly from [5, 6]
Mix coefficient $\alpha$ Randomly from [0.3, 0.7]
Mix timestep $T$ Randomly from [20, 45]
Generated images (PACS) 32k per class
Generated images (VLCS) 32k per class
Generated images (OfficeHome) 16k per class
Number of Generated Domains The impact of varying the number of generated domains on model performance was rigorously evaluated, as summarized in Table 9. This analysis aimed to understand how different combinations of augmented domains influence the overall accuracy across various dataset domains such as Art, Cartoon, Photo, and Sketch. By integrating diverse domain combinations, identified by IDs (as defined in Table 10), we observed improvement gain when we add more generated domain of different combinations. Notably, all possible combinations of augmented domains (3 new domains for PACS, VLCS and OfficeHome) were utilized as the final method, leveraging the full spectrum of available data domains.
Table 9: Analysis of the impact of utilizing different numbers/combinations of generated domains on final model performance across the PACS dataset domains. For definitions of each augmented domain (ID0, ID1, ID2), see Table 10.
Method Augmented Domains Accuracy (%) Art Cartoon Photo Sketch Avg. SWAD (reproduced) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.45 SWAD + FPS ID0 91.03 $\pm$0.5 83.87 $\pm$0.6 97.75 $\pm$0.3 85.77 $\pm$0.4 89.61 $\pm$0.30 SWAD + FPS ID1 91.01 $\pm$0.6 85.06 $\pm$1.3 97.90 $\pm$0.3 83.64 $\pm$0.4 89.40 $\pm$0.65 SWAD + FPS ID2 91.46 $\pm$0.3 85.22 $\pm$0.8 97.88 $\pm$0.2 84.27 $\pm$0.3 89.71 $\pm$0.40 SWAD + FPS ID0 + ID1 91.52 $\pm$0.0 85.87 $\pm$0.7 98.03 $\pm$0.3 85.70 $\pm$1.0 90.28 $\pm$0.50 SWAD + FPS ID1 + ID2 91.62 $\pm$0.8 85.57 $\pm$0.4 98.20 $\pm$0.3 83.88 $\pm$0.6 89.82 $\pm$0.53 SWAD + FPS ID0 + ID2 91.52 $\pm$0.1 84.54 $\pm$0.5 98.28 $\pm$0.1 86.45 $\pm$0.8 90.20 $\pm$0.38 SWAD + FPS ID0 + ID1 + ID2 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.35
Table 10: Explanation of Augmented Domains ID Definitions for Each Target Domain of PACS dataset.
Augmented Domains Art Cartoon Photo Sketch ID0 Cartoon + Photo Art + Photo Art + Cartoon Art + Cartoon ID1 Cartoon + Sketch Art + Sketch Art + Sketch Art + Photo ID2 Photo + Sketch Photo + Sketch Cartoon + Sketch Cartoon + Photo
Appendix C Detailed Results
Here are the comprehensive tables containing all the detailed information that was summarized in the main paper.
Table 11: Deatailed Comparison of FDS with the state-of-the-art on PACS dataset. numbers marked are reproduced in-house.
Method Source Domains Art Cartoon Photo Sketch Avg. CDANN 84.6 $\pm$1.8 75.5 $\pm$0.9 96.8 $\pm$0.3 73.5 $\pm$0.6 82.60 $\pm$0.9 MASF 82.9 80.5 95.0 72.3 82.7 DMG 82.6 78.1 94.5 78.3 83.4 IRM 84.8 $\pm$1.3 76.4 $\pm$1.1 96.7 $\pm$0.6 76.1 $\pm$1.0 83.50 $\pm$1.0 MetaReg 87.2 79.2 97.6 70.3 83.6 DANN 86.4 $\pm$0.8 77.4 $\pm$0.8 97.3 $\pm$0.4 73.5 $\pm$2.3 83.65 $\pm$1.1 ERM 85.7 $\pm$0.6 77.1 $\pm$0.8 97.4 $\pm$0.4 76.6 $\pm$0.7 84.20 $\pm$0.6 GroupDRO 83.5 $\pm$0.9 79.1 $\pm$0.6 96.7 $\pm$0.3 78.3 $\pm$2.0 84.40 $\pm$0.9 MTL 87.5 $\pm$0.8 77.1 $\pm$0.5 96.4 $\pm$0.8 77.3 $\pm$1.8 84.58 $\pm$0.9 I-Mixup 86.1 $\pm$0.5 78.9 $\pm$0.8 97.6 $\pm$0.1 75.8 $\pm$1.8 84.60 $\pm$0.8 MMD 86.1 $\pm$1.4 79.4 $\pm$0.9 96.6 $\pm$0.2 76.5 $\pm$0.5 84.65 $\pm$0.7 VREx 86.0 $\pm$1.6 79.1 $\pm$0.6 96.9 $\pm$0.5 77.7 $\pm$1.7 84.93 $\pm$1.1 MLDG 85.5 $\pm$1.4 80.1 $\pm$1.7 97.4 $\pm$0.3 76.6 $\pm$1.1 84.90 $\pm$1.1 ARM 86.8 $\pm$0.6 76.8 $\pm$0.5 97.4 $\pm$0.3 79.3 $\pm$1.2 85.08 $\pm$0.6 RSC 85.4 $\pm$0.8 79.7 $\pm$1.8 97.6 $\pm$0.3 78.2 $\pm$1.2 85.23 $\pm$1.0 Mixstyle 86.8 $\pm$0.5 79.0 $\pm$1.4 96.6 $\pm$0.1 78.5 $\pm$2.3 85.23 $\pm$1.1 ER 87.5 79.3 98.3 76.3 85.3 pAdaIN 85.8 81.1 97.2 77.4 85.4 EISNet 86.6 81.5 97.1 78.1 85.8 CORAL 88.3 $\pm$0.2 80.0 $\pm$0.5 97.5 $\pm$0.3 78.8 $\pm$1.3 86.15 $\pm$0.6 SagNet 87.4 $\pm$1.0 80.7 $\pm$0.6 97.1 $\pm$0.1 80.0 $\pm$0.4 86.30 $\pm$0.5 DSON 87.0 80.6 96.0 82.9 86.6 SWAD 89.3 $\pm$0.2 83.4 $\pm$0.6 97.3 $\pm$0.3 82.5 $\pm$0.5 88.13 $\pm$0.4 ERM 86.94 $\pm$0.6 80.21 $\pm$0.7 96.61 $\pm$0.4 74.45 $\pm$2.9 84.30 $\pm$1.1 ERM + FDS (ours) 86.27 $\pm$3.1 84.24 $\pm$0.7 97.08 $\pm$0.9 81.91 $\pm$1.9 87.38 $\pm$1.6 SWAD 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$0.1 88.11 $\pm$0.4 SWAD + FDS (ours) 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
Table 12: Deatailed Comparison of FDS with the state-of-the-art on VLCS dataset. numbers marked are reproduced in-house.
Method Source Domains Caltech101 LabelMe SUN09 VOC2007 Avg. GroupDRO 97.3 $\pm$0.3 63.4 $\pm$0.9 69.5 $\pm$0.8 76.7 $\pm$0.7 76.73 $\pm$0.7 RSC 97.9 $\pm$0.1 62.5 $\pm$0.7 72.3 $\pm$1.2 75.6 $\pm$0.8 77.08 $\pm$0.7 MLDG 97.4 $\pm$0.2 65.2 $\pm$0.7 71.0 $\pm$1.4 75.3 $\pm$1.0 77.23 $\pm$0.8 MTL 97.8 $\pm$0.4 64.3 $\pm$0.3 71.5 $\pm$0.7 75.3 $\pm$1.7 77.23 $\pm$0.7 ERM 98.0 $\pm$0.3 64.7 $\pm$1.2 71.4 $\pm$1.2 75.2 $\pm$1.6 77.33 $\pm$1.8 I-Mixup 98.3 $\pm$0.6 64.8 $\pm$1.0 72.1 $\pm$0.5 74.3 $\pm$0.8 77.38 $\pm$0.7 ERM 97.7 $\pm$0.4 64.3 $\pm$0.9 73.4 $\pm$0.5 74.6 $\pm$1.3 77.50 $\pm$0.8 MMD 97.7 $\pm$0.1 64.0 $\pm$1.1 72.8 $\pm$0.2 75.3 $\pm$3.3 77.45 $\pm$1.1 CDANN 97.1 $\pm$0.3 65.1 $\pm$1.2 70.7 $\pm$0.8 77.1 $\pm$1.5 77.50 $\pm$0.9 ARM 98.7 $\pm$0.2 63.6 $\pm$0.7 71.3 $\pm$1.2 76.7 $\pm$0.6 77.58 $\pm$0.6 SagNet 97.9 $\pm$0.4 64.5 $\pm$0.5 71.4 $\pm$1.3 77.5 $\pm$0.5 77.83 $\pm$0.6 Mixstyle 98.6 $\pm$0.3 64.5 $\pm$1.1 72.6 $\pm$0.5 75.7 $\pm$1.7 77.85 $\pm$0.9 VREx 98.4 $\pm$0.3 64.4 $\pm$1.4 74.1 $\pm$0.4 76.2 $\pm$1.3 78.28 $\pm$0.8 IRM 98.6 $\pm$0.1 64.9 $\pm$0.9 73.4 $\pm$0.6 77.3 $\pm$0.9 78.55 $\pm$0.6 DANN 99.0 $\pm$0.3 65.1 $\pm$1.4 73.1 $\pm$0.3 77.2 $\pm$0.6 78.60 $\pm$0.6 CORAL 98.3 $\pm$0.1 66.1 $\pm$1.2 73.4 $\pm$0.3 77.5 $\pm$1.2 78.83 $\pm$0.7 SWAD 98.8 $\pm$0.1 63.3 $\pm$0.3 75.3 $\pm$0.5 79.2 $\pm$0.6 79.15 $\pm$0.3 ERM 96.91 $\pm$1.4 64.13 $\pm$1.4 71.11 $\pm$1.5 72.78 $\pm$0.9 76.23 $\pm$1.3 ERM + FDS (ours) 99.35 $\pm$0.2 63.71 $\pm$1.7 75.16 $\pm$3.1 79.60 $\pm$1.5 79.46 $\pm$1.6 SWAD 98.65 $\pm$0.2 63.89 $\pm$0.3 74.32 $\pm$1.1 78.61 $\pm$0.6 78.87 $\pm$0.5 SWAD + FDS (ours) 99.47 $\pm$0.2 62.89 $\pm$0.2 76.94 $\pm$0. 79.60 $\pm$1.3 79.73 $\pm$0.5
Table 13: Deatailed Comparison of FDS with the state-of-the-art on OfficeHome dataset. numbers marked are reproduced in-house.
Method Source Domains Art Clipart Product RealWorld Avg. Mixstyle 51.1 $\pm$0.3 53.2 $\pm$0.4 68.2 $\pm$0.7 69.2 $\pm$0.6 60.43 $\pm$0.5 IRM 58.9 $\pm$2.3 52.2 $\pm$1.6 72.1 $\pm$2.9 74.0 $\pm$2.5 64.30 $\pm$2.3 ARM 58.9 $\pm$0.8 51.0 $\pm$0.5 74.1 $\pm$0.1 75.2 $\pm$0.3 64.80 $\pm$0.4 RSC 60.7 $\pm$1.4 51.4 $\pm$0.3 74.8 $\pm$1.1 75.1 $\pm$1.3 65.50 $\pm$1.1 CDANN 61.5 $\pm$1.4 50.4 $\pm$2.4 74.4 $\pm$0.9 76.6 $\pm$0.8 65.73 $\pm$1.3 DANN 59.9 $\pm$1.3 53.0 $\pm$0.3 73.6 $\pm$0.7 76.9 $\pm$0.5 65.85 $\pm$0.7 GroupDRO 60.4 $\pm$0.7 52.7 $\pm$1.0 75.0 $\pm$0.7 76.0 $\pm$0.7 66.03 $\pm$0.8 MMD 60.4 $\pm$0.2 53.3 $\pm$0.3 74.3 $\pm$0.1 77.4 $\pm$0.6 66.35 $\pm$0.3 MTL 61.5 $\pm$0.7 52.4 $\pm$0.6 74.9 $\pm$0.4 76.8 $\pm$0.4 66.40 $\pm$0.5 VREx 60.7 $\pm$0.9 53.0 $\pm$0.9 75.3 $\pm$0.1 76.6 $\pm$0.5 66.40 $\pm$0.6 MLDG 61.5 $\pm$0.9 53.2 $\pm$0.6 75.0 $\pm$1.2 77.5 $\pm$0.4 66.80 $\pm$0.7 ERM 63.1 $\pm$0.3 51.9 $\pm$0.4 77.2 $\pm$0.5 78.1 $\pm$0.2 67.58 $\pm$0.3 I-Mixup 62.4 $\pm$0.8 54.8 $\pm$0.6 76.9 $\pm$0.3 78.3 $\pm$0.2 68.10 $\pm$0.5 SagNet 63.4 $\pm$0.2 54.8 $\pm$0.4 75.8 $\pm$0.4 78.3 $\pm$0.3 68.08 $\pm$0.3 CORAL 65.3 $\pm$0.4 54.4 $\pm$0.5 76.5 $\pm$0.1 78.4 $\pm$0.5 68.65 $\pm$0.4 SWAD 66.1 $\pm$0.4 57.7 $\pm$0.4 78.4 $\pm$0.1 80.2 $\pm$0.2 70.60 $\pm$0.3 ERM 59.54 $\pm$2.1 51.25 $\pm$1.3 73.81 $\pm$0.8 73.78 $\pm$0.2 64.60 $\pm$1.1 ERM + FDS (ours) 60.78 $\pm$1.4 55.90 $\pm$1.6 79.54 $\pm$0.5 80.45 $\pm$0.8 69.17 $\pm$1.1 SWAD 65.86 $\pm$0.9 56.81 $\pm$0.4 78.66 $\pm$0.3 80.04 $\pm$0.2 70.34 $\pm$0.4 SWAD + FDS (ours) 67.34 $\pm$0.8 60.52 $\pm$0.5 82.56 $\pm$0.1 83.62 $\pm$0.3 73.51 $\pm$0.4
Figure 6: Synthetic images from interpolating between “art” and “photo” domains of PACS, with selected images showcasing a blend of artistic and realistic features (top two rows) and non-selected images (bottom rows) due to class mismatches and low entropy.
Figure 7: Interpolation between “art” and “sketch” in PACS highlights selected images (top rows) merging textures and outlines, and non-selected images (bottom rows) for failing selection criteria.
Figure 8: Results from “photo” and “sketch” domain interpolation in PACS, with selected synthetic images (top rows) and non-selected due to predictability and class misalignment (bottom rows).
Figure 9: Inter-domain Transition from “photo” to “art”. This sequence illustrates how varying $\alpha$ from 0.0 (purely photorealistic images) to 1.0 (purely artistic representations) enables the model to seamlessly blend photographic realism with artistic expression, demonstrating a smooth progression from real-world imagery to stylized art.
Figure 10: Inter-domain Transition from “sketch” to “art”. Displayed here is the transformation that occurs as $\alpha$ is adjusted, beginning with 0.0 (pure sketches) and moving towards 1.0 (fully art-inspired images). The model effectively infuses basic sketches with complex textures and colors, transitioning from minimalistic line art to detailed and vibrant artistic images.
Figure 11: Inter-domain Transition from “sketch” to “photo”. This figure demonstrates the capability of the model to morph sketches into photorealistic images by altering $\alpha$ from 0.0 (entirely sketch-based) to 1.0 (completely photorealistic). The transition highlights the model’s proficiency in enriching simple outlines with lifelike details and textures, bridging the gap between abstract sketches and reality.